# Tensor Gaussian Processes: Efficient Solvers for Nonlinear PDEs

### v1

@@ 5 Related Work
While PINNs have achieved many success stories, e.g.(Raissi et al.,, 2020; Jin et al.,, 2021; Sahli Costabal et al.,, 2020; Li et al.,, 2023), the training of PINN is often lengthy and challenging, partly because applying differential operators over NNs can complicate the loss landscape (Krishnapriyan et al.,, 2021). Alternatively, early works such as (Graepel,, 2003; Raissi et al.,, 2017) developed GP models to solve linear PDEs from noisy observations of the source terms. Chen et al., 2021b ; Long et al., (2022) further extended this direction by developing a kernel method capable of addressing both linear and nonlinear PDEs. Batlle et al., (2023) developed a rigorous convergence framework, establishing both convergence guarantees and rates of (Chen et al., 2021b, ). To mitigate the scalability issue for massive collocation sets, Chen et al., (2023) proposed a sparse approximation technique based on the sparse inverse Cholesky factorization (Schafer et al.,, 2021).
Xu et al., (2025) proposed using a standard kernel interpolation framework to approximate the PDE solution, and induced a Kronecker product structure to simply the computation. The efficiency of Kronecker-structured GP models has been studied in prior works (Saatcci,, 2012; Wilson and Nickisch,, 2015; Wilson et al.,, 2015; Izmailov et al.,, 2018). Fang et al., (2023) leveraged a similar idea for solving high-frequency and multiscale PDEs, using a spectral mixture kernel along each input dimension to capture dominant frequencies in the kernel function. Broader discussions on Bayesian approaches to PDEs are given in (Owhadi,, 2015).
The idea of using separable function approximations for solving PDEs was introduced by Beylkin and Mohlenkamp, (2002) and has seen rapid development in recent years. Notable works include Richter et al., (2021), Oster et al., (2022), and Fackeldey et al., (2022), which apply tensor decomposition techniques — often with random sampling — to solve PDEs and/or stochastic differential equations efficiently. A comprehensive review of deterministic methods in this context can be found in Bachmayr, (2023).
@@ 6 Numerical Experiments
We evaluated our method on five commonly-used benchmark PDE families from the literature on machine learning PDE solvers (Raissi et al., 2019a, ; Chen et al., 2021c, ; Xu et al.,, 2024): viscous Burger’s equations, nonlinear elliptic PDEs, Eikonal PDEs, Allen-Cahn equations, and nonlinear Darcy Flow equations. The details are provided in Appendix Section A. We denote our method using partial freezing as TGPS-PF and Newton’s method as TGPS-NT. We compared against two state-of-the-art GP/kernel-based solvers: (1) DAKS (Derivative-Augmented Kernel Solver) Chen et al., 2021c , which augments the GP covariance matrix with derivative information to estimate solution values and their derivatives at the collocation points. (2) SKS (Simple Kernel-based Solver) (Xu et al.,, 2025), which employs standard GP interpolation. We also compared with (3) PINN (Raissi et al., 2019a, ). SKS, DAKS and TGPS were implemented with JAX (Frostig et al.,, 2018), while PINN with PyTorch (Paszke et al.,, 2019). Hyperparameters and settings are detailed in Appendix Section B.
@@ 6.1 Solution Accuracy
Simpler Cases. We first evaluated all methods on relatively simple benchmarks, where only a modest number of collocation points is required. Specifically, we considered Burgers’ equation (17) with viscosity $\nu=0.02$, the nonlinear elliptic PDE (18), and the Eikonal PDE (19) — the same test cases as in (Chen et al., 2021b, ). Following the setups in (Chen et al., 2021b, ; Xu et al.,, 2025), the number of collocation points was varied as 600, 1200, 2400, 4800 for Burgers’ equation and 300, 600, 1200, 2400 for the nonlinear elliptic and Eikonal PDEs. DAKS employed randomly sampled collocation points, whereas SKS used a regularly spaced square grid with a comparable number of points. Notably, DAKS performed worse under gridded collocation, as also observed in (Xu et al.,, 2025). We ran PINN and our method on both randomly sampled points and regularly spaced grids, and report the best outcomes. For methods using random sampling, each experiment was repeated five times and the average relative $L^{2}$ error was recorded. As shown in Table 1, our method — both TGPS-PF and TGPS-NT — consistently achieves the highest solution accuracy. The only exception is when solving Burgers’ equation ($\nu=0.02$) with 600 collocation points, where TGPS-NT performs slightly worse than PINN but still ranks third. Note that for all these PDEs the input dimension is $d=2$, under which the TR and CP decomposition forms are equivalent (see Section 3.1). Accordingly, we do not introduce additional notation to distinguish between them.
@@
We then increased the number of collocation points to the range of 6.4K – 40K. In this regime, DAKS became prohibitively expensive, so no results are reported. As shown in Table 2, TGPS substantially reduce errors, achieving $10^{-4}$ for Burgers’ equation with 2828K–3232K collocation points, and $10^{-6}$ for the 2D Allen-Cahn equation (both $a=15$ and $a=20$) across all the cases. While SKS also improves significantly, its accuracy is consistently worse than TGPS (except for Burgers’ equation with 88K collocation points, where SKS slightly outperforms TGPS-PF), often by one order of magnitude. PINN, by contrast, only reaches a relative $L^{2}$ error of $10^{-2}$ on Burgers’ equation and remains highly inaccurate on Allen-Cahn (relative $L^{2}$ error exceeding one), with little benefit from additional collocation points. This poor performance is likely due to the relatively high-frequency components in the Allen-Cahn solution (see (20)), which neural networks struggle to capture because of their known spectral bias (Rahaman et al.,, 2019).
@@
Comparison with Conventional Numerical Methods. We compared against two established numerical methods. The first is the P2 Galerkin Finite Element Method (FEM) (Barrett and Liu,, 1993; Brenner and Scott,, 2008), and the second is a robust finite difference (FD) scheme. Details are provided in Appendix Section G. We conducted experiments on the nonlinear elliptic PDE (18) and the 2D Allen–Cahn equation (20) with $a=15$ and $a=20$, where ground-truth solutions are available for fair comparison. For FEM, we set the mesh spacing to match the collocation grids used in SKS. The FD scheme used the same grids as SKS. As reported in Appendix Table 4, our method (TGPS) consistently outperforms the conventional numerical solvers, often by several orders of magnitude in error.
@@ Appendix A PDE Benchmarks
where $\nu$ is the viscosity. The solution is computed via the Cole-Hopf transformation with quadrature (Chen et al., 2021a, ). We considered two cases: $\nu=0.02$ and $\nu=0.001$.
Nonlinear Elliptic PDE. We then tested on a nonlinear elliptic equation (Chen et al., 2021a, ),
@@
Eikonal PDE. The third is a regularized Eikonal equation (Chen et al., 2021b, ; Xu et al.,, 2025),
@@
where $\Omega=[0,1]^{2}$, $g({\bf x})=1$, and $\epsilon=0.1$. The solution was calculated from a highly-resolved finite difference solver provided by (Chen et al., 2021b, ).
Allen-Cahn Equation. Fourth, we considered a stationary Allen-Cahn equation with Dirichlet boundary conditions, generalized from the benchmark used in (Xu et al.,, 2025),
@@
Nonlinear Darcy Flow. Fifth, we employed a 6D nonlinear Darcy flow equation with Dirichlet boundary conditions (Batlle et al.,, 2025):
@@
where each $x_{i}\in[0,1]$, $c(x_{1},\cdots,x_{6})=\exp{(\sin(\sum\nolimits_{i=1}^{6}\cos\left(x_{i}\right)))}$, the solution is crafted as $u=\exp{\left(\sin\left(\beta\sum\nolimits_{i=1}^{6}\cos\left(x_{j}\right)\right)\right)}$, and $a$ is computed based on the PDE. We set $\beta=6$, which is more challenging than the case used in (Batlle et al.,, 2025).
@@ Appendix B Method Details
  • SKS. We used the original JAX implementation [^1] https://github.com/BayesianAIGroup/Efficient-Kernel-PDE-Solver. The training is conduced via ADAM optimization with initial learning rate $10^{-3}$. The maximum number of iterations was set to 1M. In the training process, SKS does not sample mini-batches of collocation points to compute stochastic gradients. Instead, the full gradient is computed from the training objective at each step, and then fed into ADAM optimizer to update the momentum online and to adjust element-wise step-size. The optimization is stopped if the training objective does not improve for 1K updates. SKS employed Square Exponential (SE) kernel with different length-scales across the input dimensions. Alternative kernels, such as the Matérn kernel, led to inferior performance. The length-scale hyperparameters were selected from a grid search, as detailed in the original paper (Xu et al.,, 2025). The nugget term was selected from $\{5\times 10^{-5},10^{-5},5\times 10^{-6},10^{-6},\ldots,10^{-13}\}$.
  • DAKS. We used the original JAX implementation provided by the authors [^2] https://github.com/yifanc96/NonLinPDEs-GPsolver. The training is performed using relaxed Gauss-Newton optimization. Note that applying the same method to train SKS almost always led to divergence. The kernel was selected from among the squared exponential (SE) kernel and the Matérn kernel with degrees of freedom 3/2 or 5/2. The nugget term was selected from the set $\{5\times 10^{-5},10^{-5},5\times 10^{-6},10^{-6},\ldots,10^{-13}\}$. Hyperparameters were chosen following the same procedure as for SKS. For solving the nonlinear elliptic PDE with DAKS, however, we employed its default strategy (Chen et al., 2021b, ), which adaptively assigns nugget values to the two sub-blocks of the Gram matrix; this yielded the best performance for DAKS.
@@
The proof of Lemma 4.1 is based on the existing results for the Tucker decomposition in (Griebel and Harbrecht,, 2023). In particular, (Griebel and Harbrecht,, 2023, Theorem 2) states that for the Tucker format,
@@
The proof of Lemma 4.2 builds on (Griebel and Harbrecht,, 2023, Theorem 3). In particular, the approximation results apply to the TT format (Oseledets,, 2011) by invoking (Griebel and Harbrecht,, 2023, Theorems 4-5 and Remark 2), where we specialize to the case in which each factor function has input dimension one. Since the TT format is a special case of the TR format with $R_{0}=R_{d}=1$, these approximation results carry over to the TR format as well. ∎
@@ Appendix D Bound of RKHS Norm
where ${\hat{f}}^{i}_{r}$ is from the approximation $^{\widehat{u}}$ in (36). This is a standard kernel regression problem. Let us denote the minimizer of (39) by ${\hat{f}}^{i}_{r{\mathcal{M}}}$. According to the optimal recovery theorem (Owhadi and Scovel,, 2019), we have ${\hat{f}}^{i}_{r{\mathcal{M}}}$ takes the kernel interpolation form (6), and
@@
Obviously, ${\widehat{u}}_{\mathcal{M}}({\bf x}_{m})-{\widehat{u}}({\bf x}_{m})=0$ for every ${\bf x}_{m}\in{\mathcal{M}}_{\Omega}$. According to the sampling inequality ( Proposition A.1 of (Batlle et al.,, 2023)), when $h$ is sufficiently small (note that the fill-in distance $h_{\Omega}\leq h$),
@@
Step 3. Let us denote by $u^{\dagger}$ the solution of problem (7). We then analyze the error of $u^{\dagger}$. Using an idea similar to (Xu et al.,, 2025), we define two error functions,

### v2

@@ 5 Related Work
While PINNs have achieved many success stories, e.g.(Raissi et al., 2020; Jin et al., 2021; Sahli Costabal et al., 2020; Li et al., 2023), the training of PINN is often lengthy and challenging, partly because applying differential operators over NNs can complicate the loss landscape (Krishnapriyan et al., 2021). Alternatively, early works such as (Graepel, 2003; Raissi et al., 2017) developed GP models to solve linear PDEs from noisy observations of the source terms. Chen et al. (2021b); Long et al. (2022) further extended this direction by developing a kernel method capable of addressing both linear and nonlinear PDEs. Batlle et al. (2023) developed a rigorous convergence framework, establishing both convergence guarantees and rates of (Chen et al., 2021b). To mitigate the scalability issue for massive collocation sets, Chen et al. (2023) proposed a sparse approximation technique based on the sparse inverse Cholesky factorization (Schafer et al., 2021).
Xu et al. (2025) proposed using a standard kernel interpolation framework to approximate the PDE solution, and induced a Kronecker product structure to simply the computation. The efficiency of Kronecker-structured GP models has been studied in prior works (Saatçi, 2012; Wilson and Nickisch, 2015; Wilson et al., 2015; Izmailov et al., 2018). Fang et al. (2023) leveraged a similar idea for solving high-frequency and multiscale PDEs, using a spectral mixture kernel along each input dimension to capture dominant frequencies in the kernel function. Broader discussions on Bayesian approaches to PDEs are given in (Owhadi, 2015).
The idea of using separable function approximations for solving PDEs was introduced by Beylkin and Mohlenkamp (2002) and has seen rapid development in recent years. Notable works include Richter et al. (2021), Oster et al. (2022), and Fackeldey et al. (2022), which apply tensor decomposition techniques — often with random sampling — to solve PDEs and/or stochastic differential equations efficiently. A comprehensive review of deterministic methods in this context can be found in Bachmayr (2023).
@@ 6 Numerical Experiments
We evaluated our method on five commonly-used benchmark PDE families from the literature on machine learning PDE solvers (Raissi et al., 2019a; Chen et al., 2021c; Xu et al., 2024): viscous Burger’s equations, nonlinear elliptic PDEs, Eikonal PDEs, Allen-Cahn equations, and nonlinear Darcy Flow equations. The details are provided in Appendix Section A. We denote our method using partial freezing as TGPS-PF and Newton’s method as TGPS-NT. We compared against two state-of-the-art GP/kernel-based solvers: (1) DAKS (Derivative-Augmented Kernel Solver) Chen et al. (2021c), which augments the GP covariance matrix with derivative information to estimate solution values and their derivatives at the collocation points. (2) SKS (Simple Kernel-based Solver) (Xu et al., 2025), which employs standard GP interpolation. We also compared with (3) PINN (Raissi et al., 2019a). SKS, DAKS and TGPS were implemented with JAX (Frostig et al., 2018), while PINN with PyTorch (Paszke et al., 2019). Hyperparameters and settings are detailed in Appendix Section B.
@@ 6.1 Solution Accuracy
Simpler Cases. We first evaluated all methods on relatively simple benchmarks, where only a modest number of collocation points is required. Specifically, we considered Burgers’ equation (17) with viscosity $\nu=0.02$, the nonlinear elliptic PDE (18), and the Eikonal PDE (19) — the same test cases as in (Chen et al., 2021b). Following the setups in (Chen et al., 2021b; Xu et al., 2025), the number of collocation points was varied as 600, 1200, 2400, 4800 for Burgers’ equation and 300, 600, 1200, 2400 for the nonlinear elliptic and Eikonal PDEs. DAKS employed randomly sampled collocation points, whereas SKS used a regularly spaced square grid with a comparable number of points. Notably, DAKS performed worse under gridded collocation, as also observed in (Xu et al., 2025). We ran PINN and our method on both randomly sampled points and regularly spaced grids, and report the best outcomes. For methods using random sampling, each experiment was repeated five times and the average relative $L^{2}$ error was recorded. As shown in Table 1, our method — both TGPS-PF and TGPS-NT — consistently achieves the highest solution accuracy. The only exception is when solving Burgers’ equation ($\nu=0.02$) with 600 collocation points, where TGPS-NT performs slightly worse than PINN but still ranks third. Note that for all these PDEs the input dimension is $d=2$, under which the TR and CP decomposition forms are equivalent (see Section 3.1). Accordingly, we do not introduce additional notation to distinguish between them.
@@
We then increased the number of collocation points to the range of 6.4K – 40K. In this regime, DAKS became prohibitively expensive, so no results are reported. As shown in Table 2, TGPS substantially reduce errors, achieving $10^{-4}$ for Burgers’ equation with 2828K–3232K collocation points, and $10^{-6}$ for the 2D Allen-Cahn equation (both $a=15$ and $a=20$) across all the cases. While SKS also improves significantly, its accuracy is consistently worse than TGPS (except for Burgers’ equation with 88K collocation points, where SKS slightly outperforms TGPS-PF), often by one order of magnitude. PINN, by contrast, only reaches a relative $L^{2}$ error of $10^{-2}$ on Burgers’ equation and remains highly inaccurate on Allen-Cahn (relative $L^{2}$ error exceeding one), with little benefit from additional collocation points. This poor performance is likely due to the relatively high-frequency components in the Allen-Cahn solution (see (20)), which neural networks struggle to capture because of their known spectral bias (Rahaman et al., 2019).
@@
Comparison with Conventional Numerical Methods. We compared against two established numerical methods. The first is the P2 Galerkin Finite Element Method (FEM) (Barrett and Liu, 1993; Brenner and Scott, 2008), and the second is a robust finite difference (FD) scheme. Details are provided in Appendix Section G. We conducted experiments on the nonlinear elliptic PDE (18) and the 2D Allen–Cahn equation (20) with $a=15$ and $a=20$, where ground-truth solutions are available for fair comparison. For FEM, we set the mesh spacing to match the collocation grids used in SKS. The FD scheme used the same grids as SKS. As reported in Appendix Table 4, our method (TGPS) consistently outperforms the conventional numerical solvers, often by several orders of magnitude in error.
@@ 7 Conclusion
Acknowledgments
HO and SZ acknowledge support from the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation). HO acknowledges support from the Air Force Office of Scientific Research under MURI award number FOA-AFRL-AFOSR-2023-0004 (Mathematics of Digital Twins), the Department of Energy under award number DE-SC0023163 (SEA-CROGS: Scalable, Efficient, and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems), the National Science Foundations under award number 2425909 (Discovering the Law of Stress Transfer and Earthquake Dynamics in a Fault Network using a Computational Graph Discovery Approach), and from the DoD Vannevar Bush Faculty Fellowship Program under ONR award number N00014-18-1-2363. SZ acknolwedges support from NSF CAREER Award IIS-2046295, NSF OAC-2311685 (Elements: A Convergent Physics-based and Data-driven Computing Platform for Building Modeling), NSF DMS-2529112 (Collaborative Research: MATH-DT: Computationally efficient hypercomplex variable-based sensitivity methods for rapid Digital Twin model updating), and DARPA SURGE HR0011-25-C-0036 (Rapid certification of additive manufactured components using ML/AI and physics-based modeling).
@@ Appendix A PDE Benchmarks
where $\nu$ is the viscosity. The solution is computed via the Cole-Hopf transformation with quadrature (Chen et al., 2021a). We considered two cases: $\nu=0.02$ and $\nu=0.001$.
Nonlinear Elliptic PDE. We then tested on a nonlinear elliptic equation (Chen et al., 2021a),
@@
Eikonal PDE. The third is a regularized Eikonal equation (Chen et al., 2021b; Xu et al., 2025),
@@
where $\Omega=[0,1]^{2}$, $g({\bf x})=1$, and $\epsilon=0.1$. The solution was calculated from a highly-resolved finite difference solver provided by (Chen et al., 2021b).
Allen-Cahn Equation. Fourth, we considered a stationary Allen-Cahn equation with Dirichlet boundary conditions, generalized from the benchmark used in (Xu et al., 2025),
@@
Nonlinear Darcy Flow. Fifth, we employed a 6D nonlinear Darcy flow equation with Dirichlet boundary conditions (Batlle et al., 2025):
@@
where each $x_{i}\in[0,1]$, $c(x_{1},\cdots,x_{6})=\exp{(\sin(\sum\nolimits_{i=1}^{6}\cos\left(x_{i}\right)))}$, the solution is crafted as $u=\exp{\left(\sin\left(\beta\sum\nolimits_{i=1}^{6}\cos\left(x_{j}\right)\right)\right)}$, and $a$ is computed based on the PDE. We set $\beta=6$, which is more challenging than the case used in (Batlle et al., 2025).
@@ Appendix B Method Details
  • SKS. We used the original JAX implementation [^2] https://github.com/BayesianAIGroup/Efficient-Kernel-PDE-Solver. The training is conduced via ADAM optimization with initial learning rate $10^{-3}$. The maximum number of iterations was set to 1M. In the training process, SKS does not sample mini-batches of collocation points to compute stochastic gradients. Instead, the full gradient is computed from the training objective at each step, and then fed into ADAM optimizer to update the momentum online and to adjust element-wise step-size. The optimization is stopped if the training objective does not improve for 1K updates. SKS employed Square Exponential (SE) kernel with different length-scales across the input dimensions. Alternative kernels, such as the Matérn kernel, led to inferior performance. The length-scale hyperparameters were selected from a grid search, as detailed in the original paper (Xu et al., 2025). The nugget term was selected from $\{5\times 10^{-5},10^{-5},5\times 10^{-6},10^{-6},\ldots,10^{-13}\}$.
  • DAKS. We used the original JAX implementation provided by the authors [^3] https://github.com/yifanc96/NonLinPDEs-GPsolver. The training is performed using relaxed Gauss-Newton optimization. Note that applying the same method to train SKS almost always led to divergence. The kernel was selected from among the squared exponential (SE) kernel and the Matérn kernel with degrees of freedom 3/2 or 5/2. The nugget term was selected from the set $\{5\times 10^{-5},10^{-5},5\times 10^{-6},10^{-6},\ldots,10^{-13}\}$. Hyperparameters were chosen following the same procedure as for SKS. For solving the nonlinear elliptic PDE with DAKS, however, we employed its default strategy (Chen et al., 2021b), which adaptively assigns nugget values to the two sub-blocks of the Gram matrix; this yielded the best performance for DAKS.
@@
The proof of Lemma 4.1 is based on the existing results for the Tucker decomposition in (Griebel and Harbrecht, 2023). In particular, (Griebel and Harbrecht, 2023, Theorem 2) states that for the Tucker format,
@@
The proof of Lemma 4.2 builds on (Griebel and Harbrecht, 2023, Theorem 3). In particular, the approximation results apply to the TT format (Oseledets, 2011) by invoking (Griebel and Harbrecht, 2023, Theorems 4-5 and Remark 2), where we specialize to the case in which each factor function has input dimension one. Since the TT format is a special case of the TR format with $R_{0}=R_{d}=1$, these approximation results carry over to the TR format as well. ∎
@@ Appendix D Bound of RKHS Norm
where ${\hat{f}}^{i}_{r}$ is from the approximation $^{\widehat{u}}$ in (36). This is a standard kernel regression problem. Let us denote the minimizer of (39) by ${\hat{f}}^{i}_{r{\mathcal{M}}}$. According to the optimal recovery theorem (Owhadi and Scovel, 2019), we have ${\hat{f}}^{i}_{r{\mathcal{M}}}$ takes the kernel interpolation form (6), and
@@
Obviously, ${\widehat{u}}_{\mathcal{M}}({\bf x}_{m})-{\widehat{u}}({\bf x}_{m})=0$ for every ${\bf x}_{m}\in{\mathcal{M}}_{\Omega}$. According to the sampling inequality ( Proposition A.1 of (Batlle et al., 2023)), when $h$ is sufficiently small (note that the fill-in distance $h_{\Omega}\leq h$),
@@
Step 3. Let us denote by $u^{\dagger}$ the solution of problem (7). We then analyze the error of $u^{\dagger}$. Using an idea similar to (Xu et al., 2025), we define two error functions,
@@ Appendix I Irregular Domains
From a theoretical standpoint, the kernel and its hyperparameters determine the GP prior and thus the associated RKHS. The closer this RKHS matches the regularity class and characteristic scales of the true PDE solution, the better the approximation quality and stability. Hyperparameter tuning can therefore be viewed as selecting an RKHS whose inductive bias is well aligned with the target solution. While there is no universal rule for kernel and hyperparameter selection, we have found several useful heuristics that guide kernel and length-scale choices:
  • Smooth solutions (e.g., moderate-viscosity Burgers’, elliptic PDE): Gaussian kernels generally perform well.
  • Solutions with sharp gradients or low regularity (e.g., near-shocks in Burgers’ with small viscosity): less smooth kernels such as Matérn-2/3 are often more appropriate.
  • Higher-frequency structure: Smaller length-scales help capture oscillatory behavior.
For instance, in Burgers’ equation, viscosity $0.02$ was well modeled using a Gaussian kernel in space, whereas viscosity $0.001$ required switching to Matérn-2/3. For the nonlinear elliptic PDE and Allen-Cahn equation, Gaussian kernels remained effective but required different length-scales (e.g., $0.1$ for the elliptic PDE vs. $0.04$ for Allen–Cahn with $a=15$ ), consistent with their different effective frequency content.